OpenStudy (anonymous):

Can someone help me solve this problem: t=100^3 / 215 ( the 100^3/215 is in a radical)????

OpenStudy (saifoo.khan):

\[\Huge \sqrt{\frac{100^3}{215}}\]Like that?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

square root of 100 is 10 10 cubed is 1000 i am going to bet that the denominator is 225, not 215

OpenStudy (anonymous):

@upsilon you computed \[\sqrt{10}^3\] but i think it is \[\sqrt{100}^3\]

OpenStudy (anonymous):

@tammieh check your denominator and make sure it is correct. because 215 has no rational root, but the square root of 225 is 15

OpenStudy (anonymous):

so i will guess that you are looking for \[\sqrt{\frac{100^3}{225}}\]

OpenStudy (anonymous):

in which case the answer is \[\frac{1000}{15}=\frac{200}{3}\] but i could be wrong

OpenStudy (anonymous):

Okay i totally dont mean to intrude and please finish with this question before you even pay attention to me! but i was hoping if you, satellite, could look at my question real quick cause i am in dire need of some help :/ and i know you are very good at math!

OpenStudy (anonymous):

My denominator is 215, so it will stay in the radical?